the force on a charge moving with velocity (v) in a magnetic field (B) half of a maximum force.

Calculate the angle between V and B. F=qvB


1
Expert's answer
2022-03-09T15:54:54-0500

The Lorentz force is given by

"F=qvB\sin\theta"

Hence, the angle between field and velocity

"\theta=\sin^{-1}\left(\frac{F}{qvB}\right)"